# Particle motion and thermal effects around a Kalb–Ramond black hole

### v1

@@ I Introduction
Departures from conventional Lorentz invariance are often examined through the framework of the Standard Model Extension, especially in scenarios involving spontaneous symmetry breaking liu2024shadow ; 12 ; KhodadiPoDU2023 ; AraujoFilho:2024ykw ; 13 ; 11 ; 9 ; 10 ; filho2023vacuum . Among the various models explored, those incorporating a vector field—commonly referred to as the bumblebee field—have gained attention through the years. In these setups, the vector field acquires a nonzero vacuum expectation value, effectively selecting a preferred direction in spacetime and thereby distorting the isotropy of physical processes amarilo2024gravitational ; schreck2014quantum ; schreck2014quantum2 . This mechanism introduces anisotropies that alter the symmetry properties of local geometries and influence particle dynamics. Over the past years, these features have been increasingly analyzed in relation to thermodynamic effects under gravitational fields, revealing notable shifts in quantities such as entropy, temperature profiles, and specific heat across different spacetime regimes aa2021lorentz ; paperrainbow ; araujo2021higher ; araujo2021thermodynamic ; araujo2022does ; anacleto2018lorentz ; reis2021thermal ; araujo2022thermal .
@@ III.1 Index of refraction
The behavior of this quantity is illustrated in Fig. 1. As evident from the plot, the index of refraction $n(r)$ exhibits a singularity precisely at the event horizon $r_{h}$. Furthermore, Fig. 1 highlights two notable features: first, $n(r)$ remains positive within the horizon ($r<r_{h}$) and becomes negative outside it; second, as $r$ increases towards infinity, the magnitude of $n(r)$ does not vanish but instead approaches an asymptotic value given by $\ell-1$, which depends only on the parameter $\ell$.
@@ III.2 Group velocity
Fig. 3 displays this limiting behavior. As observed, $\lim\limits_{m\to 0}v_{g}(r)$ approaches the constant value 1/($1/(1-\ell)$ as $r$ increases ($r\to\infty$). Similar to the massive case, this asymptotic behavior once again highlights that the spacetime under consideration is, therefore, not asymptotically flat.
@@ III.3 Time–delay
[…]
@@ IV Interparticle potential
Moreover, having obtained the interparticle potential, it would be natural to explore another related aspect: the scattering amplitude for massive particles. However, using the Born approximation, given by $f(\gamma)=-\frac{2m}{\hbar^{2}}\frac{1}{\gamma}\int^{\infty}_{0}V(r)\,r\,\sin(\gamma r)\,\mathrm{d}r,$ we found that the integral does not converge. A similar issue arises in the massless case, where the scattering amplitude is approximated by $f(k)\;\approx\;-\,\frac{1}{4\pi}\int e^{-i\,k\cdot r}\,V(r)\,\mathrm{d}^{3}r.$ Here too, the integral fails to converge. For these reasons, a more detailed investigation of the scattering amplitude will not be pursued in the present work.
@@ V Thermodynamics from ensemble approach
To ensure conceptual clarity, the thermodynamic quantities examined in this work are formulated with respect to the natural units. As a starting point for the analysis, we introduce the classical partition function in its general formulation, following the treatment outlined in greiner2012thermodynamics
@@
The upcoming analysis will be restricted to the massless regime ($m\to 0$), which allows for fully analytical treatment of the thermodynamic quantities. Both bosonic and fermionic massless particles admit closed–form expressions; however, since their partition functions differ only by a constant factor—namely, $7/360$—we concentrate solely on bosonic modes, following the methodology adopted in Refs. filho2025modified ; araujo2023thermodynamical . The primary thermodynamic observables under consideration include the pressure $P(r,\ell)$, the internal energy $U(r,\ell)$, the entropy $S(r,\ell)$, and the heat capacity $C_{V}(r,\ell)$. It is important to mention that $F(r,\ell)=-P(r,\ell)$ represents the Helmholtz free energy.
@@ V.1 Pressure
Indeed, Eq. (28) admits multiple facts to be analyzed. We begin by examining its behavior as a function of $r$ and $\ell$, as illustrated in Fig. 7. In the left panel, the pressure $P(r,\ell)$ exhibits both positive and negative values depending on the parameters $r$ and $\ell$, suggesting the occurrence of possible phase transitions. Notably, the event horizon acts as a natural boundary that constrains such transitions. Precisely at the horizon, the pressure becomes singular, and in its immediate vicinity, $P(r,\ell)$ grows indefinitely. For the sake of comparison, the Schwarzschild case is also included. Although the plots appear to approach zero at large distances, this is merely a consequence of the chosen vertical scale, which spans the huge interval $[-10^{9},10^{9}]$.
In contrast, the right panel clearly reveals that the pressure does not vanish at spatial infinity. Instead, it asymptotically approaches a finite value, namely $\pi^{4}(1-\ell)^{3}/(45\beta^{4})$. This plot also confirms that positive pressure values arise only outside the black hole, i.e., for $r>r_{h}$
@@ V.1.1 Very close to the event horizon
Figure 8: The pressure $P(T,\ell)$ is plotted as a function of temperature $T$ in the region very close to the event horizon, i.e., at $r=1.001\times r_{h}$.
@@ V.1.2 Close to the event horizon
Figure 9: The pressure $P(T,\ell)$ is plotted as a function of temperature $T$ in the region close to the event horizon, i.e., at the photon sphere.
@@ V.1.3 Asymptotically far
Figure 10: The pressure $P(T,\ell)$ is plotted as a function of temperature $T$ in the asymptotically far region.
@@ V.2 Mean Energy
Despite appearances in the left panel suggesting that $U(r,\ell)$ tends to zero at large $r$, this visual effect stems from the extensive vertical scale selected for the plot, which covers the range $[-10^{9},10^{9}]$. Meanwhile, the right panel demonstrates that $U(r,\ell)$ remains nonzero even at spatial infinity. Rather than vanishing, it converges to the constant value $\pi^{4}(1-\ell)^{3}/(15\beta^{4})$. It is also evident that only regions outside the event horizon ($r>r_{h}$) allow for positive mean energy values.
@@ V.2.1 Very close to the event horizon
Figure 12: The mean energy $U(T,\ell)$ is plotted as a function of temperature $T$ in the region very close to the event horizon, i.e., at $r=1.001\times r_{h}$.
@@ V.2.2 Close to the event horizon
Figure 13: The mean energy $U(T,\ell)$ is plotted as a function of temperature $T$ in the region close to the event horizon, i.e., at the photon sphere.
@@ V.2.3 Asymptotically far
Figure 14: The mean energy $U(T,\ell)$ is plotted as a function of temperature $T$ in the asymptotically far region.
@@ V.3 Entropy
Here, we notice that Eq. (36) presents several noteworthy features deserving careful examination. The analysis begins by investigating how the entropy varies with respect to $r$ and $\ell$, with the results depicted in Fig. 15. As seen in the left panel, $S(r,\ell)$ can assume both positive and negative values, depending on the values ascribed to the radial coordinate and the Lorentz–violating parameter, pointing toward the potential emergence of phase transitions. Interestingly, the event horizon serves as a critical barrier beyond which such transitions are constrained. At $r=r_{h}$, the entropy diverges (as we noticed to the previous thermodynamic functions analyzed here), and just outside this radius, it increases rapidly. To provide a reference point, the Schwarzschild scenario is included. Although the curves seem to converge toward zero at large radii, this impression stems from the wide vertical range chosen for the plot, which spans from $-10^{9}$ to $10^{9}$.
@@ V.3.1 Very close to the event horizon
Figure 16: The entropy $S(T,\ell)$ is plotted as a function of temperature $T$ in the region very close to the event horizon, i.e., at $r=1.001\times r_{h}$.
@@ V.3.2 Close to the event horizon
Figure 17: The entropy $S(T,\ell)$ is plotted as a function of temperature $T$ in the region close to the event horizon, i.e., at the photon sphere.
@@ V.3.3 Asymptotically far
Figure 18: The entropy $S(T,\ell)$ is plotted as a function of temperature $T$ in the asymptotically far region.
@@ V.4 Heat capacity
Eq. (40) possesses several aspects worthy of detailed exploration. The analysis initiates by investigating its dependence on the variables $r$ and $\ell$, as shown in Fig. 19. The left panel highlights that the heat capacity $C_{V}(r,\ell)$ assumes both positive and negative values across different regimes of $r$ and $\ell$, indicating the potential manifestation of phase transitions. A crucial feature emerges at the event horizon, which naturally imposes a limit on such transitions: exactly at this boundary, $C_{V}(r,\ell)$ diverges, while in its immediate proximity, it exhibits an unbounded increase. For reference, the behavior corresponding to the Schwarzschild solution is also plotted. As happened in the previous thermodynamics functions, it is important to point out that although the graphs appear to converge toward zero at large distances, this impression results solely from the extensive vertical range adopted, spanning from $-10^{9}$ to $10^{9}$.
@@ V.4.1 Very close to the event horizon
Figure 20: The heat capacity $C_{V}(T,\ell)$ is plotted as a function of temperature $T$ in the region very close to the event horizon, i.e., at $r=1.001\times r_{h}$.
@@ V.4.2 Close to the event horizon
Figure 21: The heat capacity $C_{V}(T,\ell)$ is plotted as a function of temperature $T$ in the region close to the event horizon, i.e., at the photon sphere.
@@ V.4.3 Asymptotically far
Figure 22: The heat capacity $C_{V}(T,\ell)$ is plotted as a function of temperature $T$ in the asymptotically far region.

### v2

@@ I Introduction
Departures from conventional Lorentz invariance are often examined through the framework of the Standard Model Extension, especially in scenarios involving spontaneous symmetry breaking liu2024shadow ; 12 ; KhodadiPoDU2023 ; AraujoFilho:2024ykw ; 13 ; 11 ; 9 ; 10 ; filho2023vacuum . Among the various models explored, those incorporating a vector field—commonly referred to as the bumblebee field—have gained attention through the years. In these setups, the vector field acquires a nonzero vacuum expectation value, effectively selecting a preferred direction in spacetime and thereby distorting the isotropy of physical processes amarilo2024gravitational ; schreck2014quantum ; schreck2014quantum2 . This mechanism introduces anisotropies that alter the symmetry properties of local geometries and influence particle dynamics. Over the past years, these features have been increasingly analyzed in relation to thermodynamic effects under gravitational fields, revealing notable shifts in quantities such as entropy, temperature profiles, and specific heat across different contexts such as those found in Myers–Pospelov-type anacleto2018lorentz and Podolsky electrodynamics araujo2021thermodynamic , rainbow gravity models paperrainbow , other related higher–derivative electrodynamics araujo2021higher , bouncing cosmologies araujo2022thermal , quantum gases araujo2022does , theories with higher–dimensional operators reis2021thermal , and Lorentz–violating graviton dynamics aa2021lorentz .
@@ III General remarks
This section explores the black hole configuration that arises within the framework of Kalb–Ramond gravity, aiming to examine the broader aspects outlined earlier. In recent developments, the literature has presented two separate solutions describing such black holes. For clarity in the analysis that follows, we designate the first of these as the one introduced in Ref. yang2023static
$\begin{split}\mathrm{d}s^{2}=&-\left(\frac{1}{1-\ell}-\frac{2M}{r}\right)\mathrm{d}t^{2}+\frac{\mathrm{d}r^{2}}{\frac{1}{1-\ell}-\frac{2M}{r}}\\ &+r^{2}\mathrm{d}\theta^{2}+r^{2}\sin^{2}\mathrm{d}\varphi^{2},\end{split}$ (9)
and the second as Liu:2024oas
$\begin{split}\mathrm{d}s^{2}=&-\left(1-\frac{2M}{r}\right)\mathrm{d}t^{2}+\frac{(1-\ell)}{1-\frac{2M}{r}}\,\mathrm{d}r^{2}+r^{2}\mathrm{d}\theta^{2}\\ &+r^{2}\sin^{2}\mathrm{d}\varphi^{2}.\end{split}$ (10)
The parameter $\ell$ originates from the coupling between the gravitational sector and an antisymmetric tensor field, defined as $\ell=\xi_{2}b^{2}/2$, where $\xi_{2}$ is a dimensionless coupling coefficient and $b^{2}=b_{\mu\nu}b^{\mu\nu}$ represents the norm of the vacuum expectation value associated with the field yang2023static .
Among the black hole solutions proposed within this context, one particular configuration has been subject to a wide range of analyses. The dynamics of Vlasov–type matter accreting onto the black hole have been investigated in jiang2024accretion , while jumaniyozov2024circular explored the structure of circular geodesics and their relevance for quasi–periodic oscillations. Constraints arising from spontaneous Lorentz symmetry breaking were examined in junior2024spontaneous . Additional studies addressed the deflection of light in strong–field lensing scenarios junior2024gravitational , the spectral distribution of emitted particles through greybody factors guo2024quasinormal , particle creation araujo2025particleasda , and the response of the geometry under perturbations via quasinormal mode analysis araujo2024exploring . The model was later generalized to include electric charge duan2024electrically , prompting further phenomenological investigations into thermal properties, radiation spectra, and shadow formation hosseinifar2024shadows ; Zahid:2024ohn ; aa2024antisymmetric ; al-Badawi:2024pdx ; heidari2024impact ; chen2024thermal .
A separate solution, closely related but distinguished by a sign change in $\ell$, has also appeared in the literature. This variant shares formal features with bumblebee gravity under the metric formalism and has been applied to examine entanglement degradation in Lorentz–violating backgrounds liu2024lorentz .
In the remainder of this section, we restrict our attention to the first solution introduced in Eq. (9), which forms the basis for the analysis that follows.
@@ III.1 Index of refraction
The behavior of this quantity is illustrated in Fig. 1. As evident from the plot, the index of refraction $n(r)$ exhibits a singularity precisely at the event horizon $r_{h}$. Furthermore, Fig. 1 highlights two notable features: first, $n(r)$ remains positive within the horizon ($r<r_{h}$) and becomes negative outside it; second, as $r$ increases towards infinity, the magnitude of $n(r)$ does not vanish but instead approaches an asymptotic value given by $\ell-1$, which depends only on the parameter $\ell$. Here and in all subsequent analyses, we have set the mass parameter to $M=1$. Otherwise, the radial coordinate in the plots would need to be rescaled as $r/M$ on the xx–axis.
@@ III.2 Group velocity
Fig. 3 displays this limiting behavior. As observed, $\lim\limits_{m\to 0}v_{g}(r)$ approaches the constant value $1/(1-\ell)$ as $r$ increases ($r\to\infty$). Similar to the massive case, this asymptotic behavior once again highlights that the spacetime under consideration is, therefore, not asymptotically flat.
@@ III.3 Time–delay
Now, by considering the time–delay function $\Delta t(d)\;=\;(\ell-1)\,\Bigl{[}\,2\,(\ell-1)\,M\bigl{(}\,\ln\bigl{[}d_{2}+2(\ell-1)M\bigr{]}-\ln\bigl{[}d_{1}+2(\ell-1)M\bigr{]}\bigr{)}\;+\;d_{1}\;-\;d_{2}\Bigr{]}$. We take the supermassive black hole at the center of the Milky Way (Sagittarius A*) with a mass $M_{\mathrm{SgrA^{*}}}\;\approx\;4\times 10^{6}\,M_{\odot}\;\approx\;1.2\times 10^{10}\,\mathrm{m}$. We compare signal travel times from two radial distances $d_{1}\;\approx\;9.461\times 10^{15}\,\mathrm{m}\quad(1\text{ light--year})$, $d_{2}\;\approx\;9.470\times 10^{15}\,\mathrm{m}\quad(1.001\text{ light--years})$, and choose = $\ell\;=\;1.9\times 10^{-11}$ (in agreement with the bounds estimated in Ref. yang2023static ).
@@ IV Interparticle potential
Moreover, having obtained the interparticle potential, it would be natural to explore another related aspect: the scattering amplitude for massive particles. However, using the Born approximation, given by $f(\gamma)=-\frac{2m}{\hbar^{2}}\frac{1}{\gamma}\int^{\infty}_{0}V(r)\,r\,\sin(\gamma r)\,\mathrm{d}r,$ we found that the integral does not converge. A similar issue arises in the massless case, where the scattering amplitude is approximated by $f(k)\;\approx\;-\,\frac{1}{4\pi}\int e^{-i\,k\cdot r}\,V(r)\,\mathrm{d}^{3}r.$ Here too, the integral fails to converge. For these reasons, a more detailed investigation of the scattering amplitude will not be pursued in the present work. Nevertheless, a possible way to address this issue is by employing numerical methods.
@@ V Thermodynamics from ensemble approach
To ensure conceptual clarity, the thermodynamic quantities examined in this work are formulated with respect to the natural units, i.e., $c=\hbar=k_{B}=1$. As a starting point for the analysis, we introduce the classical partition function in its general formulation, following the treatment outlined in greiner2012thermodynamics
@@
The upcoming analysis will be restricted to the massless regime ($m\to 0$), which allows for fully analytical treatment of the thermodynamic quantities. Both bosonic and fermionic massless particles admit closed–form expressions; however, since their partition functions differ only by a constant factor—namely, 7/3607/360—we concentrate solely on bosonic modes, following the methodology adopted in Refs. filho2025modified ; araujo2023thermodynamical . For instance, consider the positive sign in Eq. (27), which corresponds to the partition function for fermionic modes. When computing the thermodynamic quantities from Eq. (29), we find that, after evaluating the relevant integrals, the distinction between fermionic and bosonic cases reduces to a constant factor, specifically a difference of $7/360$. The primary thermodynamic observables under consideration include the pressure $P(r,\ell)$, the internal energy $U(r,\ell)$, the entropy $S(r,\ell)$, and the heat capacity $C_{V}(r,\ell)$. It is important to mention that $F(r,\ell)=-P(r,\ell)$ represents the Helmholtz free energy.
@@ V.1 Pressure
Indeed, Eq. (30) admits multiple facts to be analyzed. We begin by examining its behavior as a function of $r$ and $\ell$, as illustrated in Fig. 7. In the left panel, the pressure $P(r,\ell)$ exhibits both positive and negative values depending on the parameters $r$ and $\ell$, suggesting the occurrence of possible phase transitions. Notably, the event horizon acts as a natural boundary that constrains such transitions. Precisely at the horizon, the pressure becomes singular, and in its immediate vicinity, $P(r,\ell)$ grows indefinitely. For the sake of comparison, the Schwarzschild case is also included. Although the plots appear to approach zero at large distances, this is merely a consequence of the chosen vertical scale, which spans the huge interval $10^{9},10^{9}]$.
In contrast, the right panel clearly reveals that the pressure does not vanish at spatial infinity. Instead, it asymptotically approaches a finite value, namely $\pi^{4}(1-\ell)^{3}/(45\beta^{4})$. This plot also confirms that positive pressure values arise only outside the black hole, i.e., for $r>r_{h}$ .
@@ V.1.1 Very close to the event horizon
Figure 8: The pressure $P(T,\ell)$ is plotted as a function of temperature $T$ (GeV) in the region very close to the event horizon, i.e., at $r=1.001\times r_{h}$.
@@ V.1.2 Close to the event horizon
Figure 9: The pressure $P(T,\ell)$ is plotted as a function of temperature $T$ (GeV) in the region close to the event horizon, i.e., at the photon sphere.
@@ V.1.3 Asymptotically far
Figure 10: The pressure $P(T,\ell)$ is plotted as a function of temperature $T$ (GeV) in the asymptotically far region.
@@ V.2 Mean Energy
Despite appearances in the left panel suggesting that $U(r,\ell)$ tends to zero at large $r$, this visual effect stems from the extensive vertical scale selected for the plot, which covers the range $10^{9},10^{9}]$. Meanwhile, the right panel demonstrates that $U(r,\ell)$ remains nonzero even at spatial infinity. Rather than vanishing, it converges to the constant value $\pi^{4}(1-\ell)^{3}/(15\beta^{4})$. It is also evident that only regions outside the event horizon ($r>r_{h}$) allow for positive mean energy values.
@@ V.2.1 Very close to the event horizon
Figure 12: The mean energy $U(T,\ell)$ is plotted as a function of temperature $T$ (GeV) in the region very close to the event horizon, i.e., at $r=1.001\times r_{h}$.
@@ V.2.2 Close to the event horizon
Figure 13: The mean energy $U(T,\ell)$ is plotted as a function of temperature $T$ (GeV) in the region close to the event horizon, i.e., at the photon sphere.
@@ V.2.3 Asymptotically far
Figure 14: The mean energy $U(T,\ell)$ is plotted as a function of temperature $T$ (GeV) in the asymptotically far region.
@@ V.3 Entropy
Here, we notice that Eq. (38) presents several noteworthy features deserving careful examination. The analysis begins by investigating how the entropy varies with respect to $r$ and $\ell$, with the results depicted in Fig. 15. As seen in the left panel, $S(r,\ell)$ can assume both positive and negative values, depending on the values ascribed to the radial coordinate and the Lorentz–violating parameter, pointing toward the potential emergence of phase transitions. Interestingly, the event horizon serves as a critical barrier beyond which such transitions are constrained. At $r=r_{h}$, the entropy diverges (as we noticed to the previous thermodynamic functions analyzed here), and just outside this radius, it increases rapidly. To provide a reference point, the Schwarzschild scenario is included. Although the curves seem to converge toward zero at large radii, this impression stems from the wide vertical range chosen for the plot, which spans from $10^{9}$ to $10^{9}$.
@@ V.3.1 Very close to the event horizon
Figure 16: The entropy $S(T,\ell)$ is plotted as a function of temperature $T$ (GeV) in the region very close to the event horizon, i.e., at $r=1.001\times r_{h}$.
@@ V.3.2 Close to the event horizon
Figure 17: The entropy $S(T,\ell)$ is plotted as a function of temperature $T$ (GeV) in the region close to the event horizon, i.e., at the photon sphere.
@@ V.3.3 Asymptotically far
Figure 18: The entropy $S(T,\ell)$ is plotted as a function of temperature $T$ (GeV) in the asymptotically far region.
@@ V.4 Heat capacity
Eq. (42) possesses several aspects worthy of detailed exploration. The analysis initiates by investigating its dependence on the variables $r$ and $\ell$, as shown in Fig. 19. The left panel highlights that the heat capacity $C_{V}(r,\ell)$ assumes both positive and negative values across different regimes of $r$ and $\ell$, indicating the potential manifestation of phase transitions. A crucial feature emerges at the event horizon, which naturally imposes a limit on such transitions: exactly at this boundary, $C_{V}(r,\ell)$ diverges, while in its immediate proximity, it exhibits an unbounded increase. For reference, the behavior corresponding to the Schwarzschild solution is also plotted. As happened in the previous thermodynamics functions, it is important to point out that although the graphs appear to converge toward zero at large distances, this impression results solely from the extensive vertical range adopted, spanning from $10^{9}$ to $10^{9}$.
@@ V.4.1 Very close to the event horizon
Figure 20: The heat capacity $C_{V}(T,\ell)$ is plotted as a function of temperature $T$ (GeV) in the region very close to the event horizon, i.e., at $r=1.001\times r_{h}$.
@@ V.4.2 Close to the event horizon
Figure 21: The heat capacity $C_{V}(T,\ell)$ is plotted as a function of temperature $T$ (GeV) in the region close to the event horizon, i.e., at the photon sphere.
@@ V.4.3 Asymptotically far
Figure 22: The heat capacity $C_{V}(T,\ell)$ is plotted as a function of temperature $T$ (GeV) in the asymptotically far region.